A can do a piece of work in 36 days, B in 54 days and C in 72 days. All of them began together but A left 8 days and B left 12 days before the completion of the work. How many days in all did C put in till the entire work was finished?

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(48546)

A can do a piece of work in 36 days,--------A's rate is 1/36 job/day
----------------------
B in 54 days -------------------------------B's rate is 1/54 job/day
C in 72 days.-------------------------------C's rate is 1/72 job/day
------------------------------------------------------------------------

All of them began together but A left 8 days and B left 12 days before the completion of the work. How many days in all did C put in till the entire work was finished?
--------------
Equation:
Let x be the number of days required to do the job:
(x-8)(1/36)+(x-12)(1/54)+x(1/72) = 1 job
---
I'll let you solve that for "x" which is how many days C worked.
=================================================================
Cheers,
Stan H.

Answer by scott8148(5885)

let x = "total days to complete work"

(x - 8)/36 + (x - 12)/54 + x/72 = 1

multiplying by 216 (LCD) ___ 6x - 48 + 4x - 48 + 3x = 216 ___ 13x = 312

x = 24